OpenStudy (anonymous):

A garden table and a bench cost $906 combined. The garden table costs $56 more than the bench. What is the cost of the bench?

OpenStudy (anonymous):

help

Vocaloid (vocaloid):

let g = price of garden table let b = price of bench g + b = 906 (since the sum of their prices is 906) g = b + 56 (since the garden table costs 56 more than the bench) with me so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

this is the first equation g + b = 906 this is the second equation g = b + 56 we can replace "g" with "b+56" in the first equation, giving us (b+56) + b = 906 can you try solving for b now?

OpenStudy (anonymous):

idk how

Vocaloid (vocaloid):

combine like terms